OpenStudy (anonymous):

HELP, WILL REWARD!! Use the quadratic formula to solve the equation. -x^2+5x=11 a.5/2 +- sqrt13/2 b.-5/2 +- sqrt13/2 c.2/5 +- sqrt13/2 d.-2/5 +- sqrt13/2

OpenStudy (anonymous):

-x^2+5x-11=0, so i only rewrote the equation right? then you use\[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

so put in you a b and c values, a= -1, b=5, c=-11

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

This is how I remember the quadratic formula bro: A negative boy couldn't decide whether to go to a radical party. He was square so he missed out on four awesome chicks. The party was over 2 a.m. Cheers

OpenStudy (anonymous):

i still dont get it?

OpenStudy (anonymous):

were u hear that at thats how i rember it to

OpenStudy (anonymous):

you will now: ax^2+bx+c=0 remember standard form? yea now substitute the values and look for sign changes and stuff and you should be solid :) CHeers

OpenStudy (anonymous):

IwishIknewwhatyousaid girl uh uh uh